gsmith2002:

Which represents the solution(s) of the system of equations, y = –x2 + 6x + 16 and y = –4x + 37? Determine the solution set algebraically. (3, 25) (–3, 49) (3, 25) and (7, 9) (–3, 49) and (–7, 65)

Shadow:

@dude previously answered this question here: https://questioncove.com/study#/updates/5aa95846784c457c75773800 Do you just not know how to do that?

gsmith2002:

the last answer givin didnt answe one of the choices may you give the answer

Shadow:

I'm afraid it isn't our policy that we straight up give out the answer. But I can help you get there. Trust me, this problem isn't that difficult. I'll break it down for you.

Shadow:

\[y = –x^2 + 6x + 16 \] \[y = –4x + 37\] Since they are both equal to the same variable, they are both equal to each other. Thus, we can do this: \[ -4x + 37= -x^2 + 6x + 16 \] \[x^2 - 10x + 21 = 0\] \[(x - 3)( x - 7) = 0\] \[x = 3, 7\] We need to test both of these variables now. \[-4(3) + 37 = -(3)^2 + 6(3) + 16\] \[25 = 25\] Now lets test 7 \[-4(7) + 37 = (7)^2 + 6(7) + 16\] \[-28 + 37 = 49 + 42 + 16\] \[9 = 107\] This is false. Therefore x = 3, not 7.

Shadow:

Now that we have x, we can solve for y. \[y = -4x + 37\] \[y = -4(3) + 37\] \[y = -12 + 37\] \[y = 25\]

Shadow:

Let me know if you have any questions regarding the steps

gsmith2002:

i got it now thanks alot !!!

Shadow:

Oh and this was posted in the Biology section. Moving it to the Mathematics seciton.

Shadow:

No problem @gsmith2002